Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
8x2 + 2x + 4 = 0 (Points : 10)
Two different irrational solutions
Two different imaginary solutions
Exactly one rational solution
Two different rational solutions



Answer by stanbon(75887)

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
8x2 + 2x + 4 = 0
----
b^2-4ac = 2^2-4*8*4 < 0
--------------
Ans: Two different imaginary solutions
========
Cheers,
Stan H.
========